Integrate sin³x dx using reduction method
1
Expert's answer
2020-02-20T08:34:27-0500

"\int sin^3x dx=\int sin^2 x\cdot sinx dx=\\
 replace\\
cos x=t\\
d(cosx)=dt\\
-sinxdx=dt\\
sin^2x=1-cos^2x\\
=-\int(1-t^2)dt=-(t-\frac{t^3}{3}+C)=\\
=-(cosx-\frac{cos^3x}{3}+C)"